Question 33 0/5 points Table 3. 400,000 Total population Working-age population # of employed # of unemployed 290,000 110,000I got this question wrong. Basic easy to understand explanation or a step-by-step if its possible. Thank you so much!
Answer #1

Answer is Option B 44.49

*Labour force participation rate is calculated by dividing the labour force with working age population.

* Labour force consists of employed and unemployed.

Working age population= 290000

Labour force= employed + unemployed

Employed = 110000

Unemployed= 19000

Labour force= 11000+19000= 129000.

Therefore, labour force participation rate = labour force/ working age population

= 129000/290000=44.48276=44.49